OpenStudy (anonymous):

Choose the correct simplification of the expression 8 times b all over a to the power of negative 5. a to the 5th power times b all over 8 8 over a to the 5th power times b 8a5b Already simplified

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (some.random.cool.kid):

@Hero

OpenStudy (some.random.cool.kid):

how to do this?

OpenStudy (some.random.cool.kid):

@jagr2713

OpenStudy (some.random.cool.kid):

@sammixboo

jagr2713 (jagr2713):

Is it\[\frac{ 8\times b }{ a^{5} }\]?

OpenStudy (some.random.cool.kid):

A-^5

jagr2713 (jagr2713):

Oh om\[\frac{ 8\times b }{ a^{-5} }\]

OpenStudy (some.random.cool.kid):

better

OpenStudy (some.random.cool.kid):

how do i solve that?

OpenStudy (some.random.cool.kid):

wait let me show you answers for better understanding

OpenStudy (some.random.cool.kid):

he correct simplification of the expression 8 times b all over a to the power of negative 5. a to the 5th power times b all over 8 8 over a to the 5th power times b 8a5b Already simplified

OpenStudy (some.random.cool.kid):

you there @jagr2713 :}

OpenStudy (some.random.cool.kid):

@Preetha @Abhisar @Hero

OpenStudy (some.random.cool.kid):

@sleepyjess

OpenStudy (some.random.cool.kid):

@EclipsedStar